# Fault Tolerant ML: Efficient Meta-Aggregation and Synchronous Training

### v2

@@ 1 Introduction
The efficiency and robustness of these aggregators can be systematically characterized, as shown in Karimireddy et al. (2020, 2021); Allouah et al. (2023); Farhadkhani et al. (2022). These studies not only outline the worst-case strengths of various aggregators but also set benchmarks for optimal efficiency, providing a comprehensive analysis of existing algorithms. Furthermore, they introduce the concept of Meta-Aggregators—advanced mechanisms that enhance the robustness of base aggregators by combining or modifying their outputs. Meta-aggregators are devised to bolster the system’s resilience, leveraging the strengths of underlying methods to achieve greater fault tolerance. Nevertheless, the application of existing meta-aggregators, e.g. Bucketing (Karimireddy et al., 2020) and Nearest Neighbor Mixing (NNM) (Allouah et al., 2023), comes with a high computational cost, highlighting a trade-off between enhanced robustness and efficiency. Notably, the cost of these meta-aggregators is $m$ folds higher compared to the standard “average” aggregator, where $m$ is the number of workers. This significant increase in computational demand hinders their practicality in high-dimensional scenarios and in situations where the number of workers is high.
@@
  • Efficient Meta-Aggregator (CTMA): We present “Centered Trimmed Meta Aggregator (CTMA)”, a novel meta-aggregator that upgrades baseline aggregators to optimal performance levels. CTMA significantly reduces computational demands typically associated with prior meta-aggregators, aligning its efficiency with that of the conventional “average” aggregator. This development makes CTMA suitable for widespread deployment in large-scale scenarios, effectively overcoming the computational challenges of previous meta-aggregators and facilitating the implementation of robust, fault-tolerant distributed machine learning systems.
  • Advanced Gradient Estimation Technique: Building on the insight that standard stochastic gradient methods are vulnerable to Byzantine disruptions, we suggest incorporating a recent double-momentum technique Levy (2023) to enhance resilience during the training process. Unlike traditional momentum-based approaches that necessitate complex parameter tuning, our approach simplifies the process by allowing the learning rate to be determined solely by the objective’s smoothness parameter $L$. This approach not only streamlines the setup but also achieves a more substantial error reduction, further mitigating the impact of Byzantine workers and enhancing the practicality of synchronous robust training in distributed ML systems. We establish theoretical guarantees for this approach within the Stochastic-Convex-Optimization (SCO)– a fundamental framework for the design and analysis of machine learning algorithms (Shalev-Shwartz et al., 2009).
@@ 2 Setting
where $\mathcal{K}\subseteq\mathbb{R}^{d}$ is a compact convex set. Thus, the objective is an average of $|\mathcal{G}|$ functions $\{f_{i}:\mathcal{K}\mapsto\mathbb{R}\}_{i\in\mathcal{G}}$, and each such $f_{i}(\cdot)$ can be written as an expectation over losses $f_{i}(\cdot,\mathbf{z}^{(i)})$ where $\mathbf{z}^{(i)}$ is drawn from some distribution $\mathcal{D}_{i}$ which is unknown to the learner. For ease of notation, in what follows, we will not explicitly denote $\mathbb{E}_{\mathbf{z}^{(i)}\sim\mathcal{D}^{(i)}}[\cdot]$ but rather use $\mathbb{E}[\cdot]$ to denote the expectation w.r.t. all randomization.
@@
$\begin{split}\mathbb{E}\left\|(\nabla f_{i}(\mathbf{x};\mathbf{z})-\nabla f_{i}(\mathbf{x}))-(\nabla f_{i}(\mathbf{y};\mathbf{z})-\nabla f_{i}(\mathbf{y}))\right\|^{2}\\ \leq\sigma_{L}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}~{}.\end{split}$ (5)
@@ 3 Robust Aggregation and Meta-Aggregators
$(c_{\delta},\delta)$-robust. Assume we have $m$ random vectors $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\in\mathbb{R}^{d}$. Also assume we have an ”honest” subset $\mathcal{G}\subseteq[m]$, implying $\{\mathbf{x}_{i}\}_{\in\ \mathcal{G}}$ are independent of each other. Finally, assume that there exists $\delta\in[0,1/2)$ such that $|\mathcal{G}|\geq(1-\delta)m$. Moreover, assume that for any $i\in\mathcal{G}$ there exist $\rho_{i}\geq 0$ such that,
@@
where  $\bar{\mathbf{x}}_{\mathcal{G}}:=\frac{1}{|\mathcal{G}|}\sum_{i\in\mathcal{G}}{\mathbf{x}}_{i}$. Then an aggregation rule $\mathcal{A}$ is called $(c_{\delta},\delta)$-robust is for any such $m$ random vectors it outputs $\hat{\mathbf{x}}\leftarrow\mathcal{A}(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})$ such that,
@@
Note that Allouah et al. (2023) has provided a different definition, capturing a rich family of robust aggregators (see Def.2 in Allouah et al. (2023)). In Lemma C.1 we show that any aggregator satisfying their definition will also satisfy our definition above. Thus, upper bounds of $c_{\delta}$ for various aggregation rules are based on the findings in Allouah et al. (2023) and are elaborated in Table 1. Furthermore, Allouah et al. (2023) has established that the optimal value for $c_{\delta}$ is $\frac{\delta}{1-2\delta}$. This optimal rate, as demonstrated in Table 1, is effectively achieved by the CWTM aggregator when $\delta\leq 1/3$. The performance of the other aggregators in Table 1, is suboptimal.
@@ 3.1 Centered Trimmed Meta Aggregator (CTMA)
To address the suboptimal performance of $(c_{\delta},\delta)$-robust aggregators, previous works have proposed ”meta-aggregators,” which enhance the performance of baseline suboptimal aggregators. Concretely, such meta-aggregators employ a baseline $(c_{\delta},\delta)$-robust aggregators in a black box manner to yield an improved robust aggregator with a factor of $O(\delta(1+c_{\delta}))$ instead of $c_{\delta}$. Thus, even if $c_{\delta}=O(1)$ (and therefore suboptimal); then after meta-aggregation we receive an order optimal factor of $O(\delta)$. This, in turn, enables the crafting of a wide class of optimal aggregators, which can be highly beneficial in practice.
Currently, two approaches exist for meta-aggregation: the Nearest Neighbor Mixing (NNM) method (Allouah et al., 2023) and the Bucketing technique (Karimireddy et al., 2020). Both have proven to be useful in theory as well as in practice. Unfortunately, the computational complexity of NNM and Bucketing is quite high and requires $O(dm^{2})$ computations in the worst case. This is substantially higher compared to the computational cost of most robust aggregators, as shown in Table 1. This issue severely limits the practicality of existing meta-aggregators in large-scale (large $d$) and massively parallel (large $m$) scenarios.
To mend the above issue, we propose an alternative meta-aggregator in Algorithm  1, named Centered Trimmed Meta Aggregator (CTMA). CTMA achieves performance comparable to NNM and Bucketing while requiring a computational cost of $O(dm+m\log m)$, which is similar (up to a logarithmic factor) to that of the standard average aggregator.
@@
It is immediate to see that Algorithm 1 requires $O(dm+m\log m)$ computations, in addition to computing $\mathbf{x}_{0}$. Furthermore, as a post-processing approach, CTMA can be effectively combined with either NNM or Bucketing without any additional cost, offering a versatile solution in optimizing robust aggregation processes. Note that the CTMA can be applied to a wide range of optimization problems, including non-convex problems.
@@ 4.2 Synchronous Robust $\mu^{2}$-SGD
$\leq 3\mathbb{E}\left\|{\mathbf{d}^{(i)}_{t}-\nabla f_{i}(\mathbf{x}_{t})}\right\|^{2}+3\mathbb{E}\left\|{\nabla f(\mathbf{x}_{t})-\bar{\mathbf{d}}_{\mathcal{G}}}\right\|^{2}$
$+3\mathbb{E}\left\|{\nabla f_{i}(\mathbf{x}_{t})-\nabla f(\mathbf{x}_{t})}\right\|^{2}$
@@ Appendix B Synchronous $\mu^{2}$-SGD Analysis
We follow similar steps as in he proof of Theorem 4.1 of Levy (2023). First, we consider the definition of $\mathbf{x}_{t}$, which leads us to the following relationship:
@@
$\leq\frac{1}{2T}\sum_{\tau\in[T]}\alpha_{1:T}\Delta_{\tau}+\alpha_{t}G^{*}D~{}.$
Here, the first inequality employs the Young’s inequality. For the second inequality, we introduce the notation $\Delta_{t}:=f(\mathbf{x}_{t})-f(\mathbf{x}^{*})$, and we follow Lemma B.4, which relates to the smoothness of the function $f$. In this step, we also set $\alpha_{0}=0$ and utilizes the property $\alpha_{\tau}^{2}\leq 2\alpha_{1:t}$, given that $\alpha_{\tau}=\tau$. The third inequality uses the Cauchy-Schwarz inequality. The last inequality follows the assumption in Eq. (1). It uses the fact that $t\leq T$ and $\Delta_{t}\geq 0$, $\forall t$. This step also incorporates the choice of an appropriate learning rate parameter $\eta\leq 1/4LT$.
@@
$\alpha_{1:t}\Delta$ $\leq\frac{1}{2T}\sum_{\tau\in[T]}\alpha_{1:T}\Delta_{\tau}+\frac{D^{2}}{2\eta}+\alpha_{t}G^{*}D+D\sum_{\tau\in[t]}\alpha_{\tau}\|{\varepsilon_{\tau}}\|~{}.$ (9)
@@
$\mathbb{E}[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})]$ $\leq\frac{1}{T^{2}}\mathcal{B}_{T}$
$=\frac{D^{2}}{2T^{2}\eta}+\frac{G^{*}D}{T}+\frac{D}{T^{2}}\sum_{\tau\in[T]}\alpha_{\tau}\mathbb{E}\|{\varepsilon_{\tau}}\|$
$\leq\frac{D^{2}}{2T^{2}\eta}+\frac{G^{*}D}{T}+\frac{D\tilde{\sigma}}{\sqrt{T}}\sqrt{{12c_{\delta}}+\frac{4}{m}}+6D\xi\sqrt{c_{\delta}}~{}.$ (12)
@@
$\frac{D^{2}}{2T^{2}\eta}\leq\frac{D\tilde{\sigma}}{\sqrt{T}}\sqrt{12c_{\delta}+\frac{4}{m}}+6D\xi\sqrt{c_{\delta}}~{}.$
@@
$\eta$ $\geq\frac{D^{2}}{2T^{2}}\cdot\frac{1}{\frac{D\tilde{\sigma}}{\sqrt{T}}\sqrt{12c_{\delta}+\frac{4}{m}}+6D\xi\sqrt{c_{\delta}}}=\frac{D}{2\tilde{\sigma}T^{1.5}\sqrt{12c_{\delta}+\frac{4}{m}}+12\xi T^{2}}~{}.$
@@
This approach allows us to employ a broad spectrum of learning rates $\eta\in[\eta_{min},\eta_{max}]$, such that $\eta_{max}/\eta_{min}=O(\sqrt{T})$. In contrast, the standard momentum method Karimireddy et al. (2020) requires a low-range of learning rates such that $\eta_{max}/\eta_{min}=O(1)$ to achieve an order optimal bound. Specifically, for a starting point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ and $\mathbf{x}^{*}:=\arg\min_{\mathbf{x}\in\mathbb{R}^{d}}f(\mathbf{x})$, the learning rate for momentum with adaptation to our notations is given by:
@@
Let us define the function $h(\mathbf{x})=f(\mathbf{x})-f(\mathbf{x}^{*})-\langle\nabla f(\mathbf{x}^{*}),\mathbf{x}-\mathbf{x}^{*}\rangle$. Given the convexity of $f(\mathbf{x})$, we have $f(\mathbf{x})-f(\mathbf{x}^{*})\geq\langle\nabla f(\mathbf{x}^{*}),\mathbf{x}-\mathbf{x}^{*}\rangle$, leading to $h(\mathbf{x})\geq 0$. As $h(\mathbf{x}^{*})=0$, this implies that $\mathbf{x}^{*}$ is the global minimum of $h$. Applying Lemma B.6, gives us,

### v3

@@ 1 Introduction
The efficiency and robustness of these aggregators can be systematically characterized, as shown in Karimireddy et al. (2020, 2021); Allouah et al. (2023); Farhadkhani et al. (2022). These studies not only outline the worst-case strengths of various aggregators but also set benchmarks for optimal efficiency, providing a comprehensive analysis of existing algorithms. Furthermore, they introduce the concept of Meta-Aggregators—advanced mechanisms that enhance the robustness of base aggregators by combining or modifying their outputs. Meta-aggregators are devised to bolster the system’s resilience, leveraging the strengths of underlying methods to achieve greater fault tolerance. The first meta-aggregator that aimed to enhance baseline robust aggregators was Bucketing (Karimireddy et al., 2020). However, this approach relies heavily on randomization, which can result in unstable performance in practice, as shown by Allouah et al. (2023). To address this, Nearest Neighbor Mixing (NNM) (Allouah et al., 2023) introduced a more effective approach in practice that enhanced the performance of any baseline robust aggregator. Nevertheless, the application of NNM comes with high computational costs, creating a trade-off between enhanced robustness and efficiency. Notably, the computational demand for NNM is $m$ times higher compared to the standard “average” aggregator, where $m$ is the number of workers. This significant increase in computational demand limits its practicality in high-dimensional scenarios and in situations where the number of workers is large.
@@
  • Efficient Meta-Aggregator (CTMA): We present “Centered Trimmed Meta Aggregator (CTMA)”, a novel meta-aggregator that upgrades baseline aggregators to optimal performance levels. CTMA offers stable performance in practice compared to Bucketing while significantly reducing computational demands typically associated with NNM, aligning its efficiency with that of the conventional “average” aggregator. This development makes CTMA suitable for widespread deployment in large-scale scenarios, balancing practical stability with low computational demands and facilitating the implementation of robust, fault-tolerant distributed machine learning systems.
  • Advanced Gradient Estimation Technique: Building on the insight that standard stochastic gradient methods are vulnerable to Byzantine disruptions, we suggest incorporating a recent double-momentum technique (Levy, 2023) to enhance resilience during the training process. Unlike traditional momentum-based approaches that necessitate complex parameter tuning, our approach simplifies the process by allowing the learning rate to be determined solely by the objective’s smoothness parameter $L$. This approach not only streamlines the setup but also achieves a more substantial error reduction, further mitigating the impact of Byzantine workers and enhancing the practicality of synchronous robust training in distributed ML systems. We establish theoretical guarantees for this approach within the Stochastic-Convex-Optimization (SCO)– a fundamental framework for the design and analysis of machine learning algorithms (Shalev-Shwartz et al., 2009).
@@ 2 Setting
where $\mathcal{K}\subseteq\mathbb{R}^{d}$ is a compact convex set. Thus, the objective is an average of $|\mathcal{G}|$ functions $\{f_{i}:\mathcal{K}\mapsto\mathbb{R}\}_{i\in\mathcal{G}}$, and each such $f_{i}(\cdot)$ can be written as an expectation over losses $f_{i}(\cdot,\mathbf{z}^{(i)})$ where $\mathbf{z}^{(i)}$ is drawn from some distribution $\mathcal{D}^{(i)}$ which is unknown to the learner. For ease of notation, in what follows, we will not explicitly denote $\mathbb{E}_{\mathbf{z}^{(i)}\sim\mathcal{D}^{(i)}}[\cdot]$ but rather use $\mathbb{E}[\cdot]$ to denote the expectation w.r.t. all randomization.
@@
$\begin{split}\mathbb{E}\left\|(\nabla f_{i}(\mathbf{x};\mathbf{z})-\nabla f_{i}(\mathbf{x}))-(\nabla f_{i}(\mathbf{y};\mathbf{z})-\nabla f_{i}(\mathbf{y}))\right\|^{2}\leq\sigma_{L}^{2}\|\mathbf{x}-\mathbf{y}\|^{2}~{}.\end{split}$ (5)
@@ 3 Robust Aggregation and Meta-Aggregators
$(c_{\delta},\delta)$-robust. Assume we have $m$ random vectors $\mathbf{x}_{1},\ldots,\mathbf{x}_{m}\in\mathbb{R}^{d}$. Also assume we have an ”honest” subset $\mathcal{G}\subseteq[m]$, implying $\{\mathbf{x}_{i}\}_{i\in\mathcal{G}}$ are independent of each other. Finally, assume that there exists $\delta\in[0,1/2)$ such that $|\mathcal{G}|\geq(1-\delta)m$. Moreover, assume that for any $i\in\mathcal{G}$ there exist $\rho_{i}\geq 0$ such that,
@@
where  $\bar{\mathbf{x}}_{\mathcal{G}}:=\frac{1}{|\mathcal{G}|}\sum_{i\in\mathcal{G}}{\mathbf{x}}_{i}$. Then an aggregation rule $\mathcal{A}$ is called $(c_{\delta},\delta)$-robust is for any such $m$ random vectors it outputs $\hat{\mathbf{x}}\leftarrow\mathcal{A}(\mathbf{x}_{1},\ldots,\mathbf{x}_{m})$ such that,
@@
Note that Allouah et al. (2023) has provided a different definition, capturing a rich family of robust aggregators (see Def.2 in Allouah et al. (2023)). In Lemma C.1, we show that any aggregator satisfying their definition will also satisfy our definition above. Thus, upper bounds of $c_{\delta}$ for various aggregation rules are based on the findings in Allouah et al. (2023) and are elaborated in Table 1. Furthermore, Allouah et al. (2023) has established that a lower bound for $c_{\delta}$ is $\frac{\delta}{1-2\delta}$ and thus, the optimal rate of $c_{\delta}<O(\delta)$ is achieved when $\delta<1/3$. Additionally, this optimal rate is effectively achieved by the CWTM aggregator, as demonstrated in Table 1. The performance of the other aggregators in Table 1 is sub-optimal.
@@ 3.1 Centered Trimmed Meta Aggregator (CTMA)
To address the sub-optimal performance of $(c_{\delta},\delta)$-robust aggregators, previous works have proposed ”meta-aggregators,” which enhance the performance of baseline sub-optimal aggregators. Concretely, such meta-aggregators employ a baseline $(c_{\delta},\delta)$-robust aggregators in a black box manner to yield an improved robust aggregator with a factor of $O(\delta(1+c_{\delta}))$ instead of $c_{\delta}$. Thus, even if $c_{\delta}=O(1)$ (and therefore sub-optimal); then after meta-aggregation we receive an order optimal factor of $O(\delta)$. This, in turn, enables the crafting of a wide class of optimal aggregators, which can be highly beneficial in practice.
Currently, two approaches exist for meta-aggregation: Bucketing (Karimireddy et al., 2020) Nearest Neighbor Mixing (NNM) (Allouah et al., 2023). Bucketing was the first proposed meta-aggregator, but it proved to be less stable in practice due to its heavy reliance on randomization (Allouah et al., 2023). To address this, NNM was introduced as a more practical alternative; however, its computational complexity is notably high, requiring $O(dm^{2})$ computations in the worst case. This is substantially higher compared to the computational cost of many prevalent robust aggregators, as shown in Table 1, and limits its practicality in large-scale (large $d$) and massively parallel (large $m$) scenarios.
To mend the above issue, we propose an alternative meta-aggregator in Algorithm  1, named Centered Trimmed Meta Aggregator (CTMA). CTMA achieves guarantees comparable to NNM and Bucketing while requiring a computational cost of $O(dm+m\log m)$, which is similar (up to a logarithmic factor) to that of the standard average aggregator. Later in this paper, we will also demonstrate its practical effectiveness.
@@
It is immediate to see that Algorithm 1 requires $O(dm+m\log m)$ computations, in addition to computing $\mathbf{x}_{0}$. Furthermore, as a post-processing approach, CTMA can be effectively combined with either NNM or Bucketing with only an additional logarithmic factor at most, offering a versatile solution in optimizing robust aggregation processes. Note that the CTMA can be applied to a wide range of optimization problems, including non-convex problems.
@@ 4.2 Synchronous Robust $\mu^{2}$-SGD
$\leq 3\mathbb{E}\left\|{\mathbf{d}^{(i)}_{t}-\nabla f_{i}(\mathbf{x}_{t})}\right\|^{2}+3\mathbb{E}\left\|{\nabla f(\mathbf{x}_{t})-\bar{\mathbf{d}}_{\mathcal{G}}}\right\|^{2}+3\mathbb{E}\left\|{\nabla f_{i}(\mathbf{x}_{t})-\nabla f(\mathbf{x}_{t})}\right\|^{2}$
@@ Appendix B Synchronous $\mu^{2}$-SGD Analysis
We follow similar steps as in the proof of Theorem 4.1 of Levy (2023). First, we consider the definition of $\mathbf{x}_{t}$, which leads us to the following relationship:
@@
$\leq\frac{1}{2T}\sum_{\tau\in[T]}\alpha_{1:\tau}\Delta_{\tau}+\alpha_{t}G^{*}D~{}.$
Here, the first inequality employs the Young’s inequality. For the second inequality, we introduce the notation $\Delta_{t}:=f(\mathbf{x}_{t})-f(\mathbf{x}^{*})$, and we follow Lemma B.4, which relates to the smoothness of the function $f$. In this step, we also set $\alpha_{0}=0$ and utilizes the property $\alpha_{\tau}^{2}\leq 2\alpha_{1:\tau}$, given that $\alpha_{\tau}=\tau$. The third inequality uses the Cauchy-Schwarz inequality. The last inequality follows the assumption in Eq. (1). It uses the fact that $t\leq T$ and $\Delta_{t}\geq 0$, $\forall t$. This step also incorporates the choice of an appropriate learning rate parameter $\eta\leq 1/4LT$.
@@
$\alpha_{1:t}\Delta$ $\leq\frac{1}{2T}\sum_{\tau\in[T]}\alpha_{1:\tau}\Delta_{\tau}+\frac{D^{2}}{2\eta}+\alpha_{t}G^{*}D+D\sum_{\tau\in[t]}\alpha_{\tau}\|{\varepsilon_{\tau}}\|~{}.$ (9)
@@
$\mathbb{E}[f(\mathbf{x}_{T})-f(\mathbf{x}^{*})]$ $\leq\frac{2}{T^{2}}B_{T}$
$=\frac{D^{2}}{T^{2}\eta}+\frac{2G^{*}D}{T}+\frac{2D}{T^{2}}\sum_{\tau\in[T]}\alpha_{\tau}\mathbb{E}\|{\varepsilon_{\tau}}\|$
$\leq\frac{D^{2}}{T^{2}\eta}+\frac{2G^{*}D}{T}+\frac{2D\tilde{\sigma}}{\sqrt{T}}\sqrt{{12c_{\delta}}+\frac{4}{m}}+12D\xi\sqrt{c_{\delta}}~{}.$ (12)
@@
$\frac{D^{2}}{T^{2}\eta}\leq\frac{2D\tilde{\sigma}}{\sqrt{T}}\sqrt{12c_{\delta}+\frac{4}{m}}+12D\xi\sqrt{c_{\delta}}~{}.$
@@
$\eta$ $\geq\frac{D^{2}}{T^{2}}\cdot\frac{1}{\frac{2D\tilde{\sigma}}{\sqrt{T}}\sqrt{12c_{\delta}+\frac{4}{m}}+12D\xi\sqrt{c_{\delta}}}=\frac{D}{2\tilde{\sigma}T^{1.5}\sqrt{12c_{\delta}+\frac{4}{m}}+12\xi T^{2}}~{}.$
@@
This approach allows us to employ a broad spectrum of learning rates $\eta\in[\eta_{min},\eta_{max}]$, such that $\eta_{max}/\eta_{min}=O(\sqrt{T})$ when considering the dominant parts at $T$. In contrast, the standard momentum method Karimireddy et al. (2020) requires a low-range of learning rates such that $\eta_{max}/\eta_{min}=O(1)$ to achieve an order optimal bound. Specifically, for a starting point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ and $\mathbf{x}^{*}:=\arg\min_{\mathbf{x}\in\mathbb{R}^{d}}f(\mathbf{x})$, the learning rate for momentum with adaptation to our notations is given by:
@@
Let us define the function $h(\mathbf{x})=f(\mathbf{x})-f(\mathbf{x}^{*})-\langle\nabla f(\mathbf{x}^{*}),\mathbf{x}-\mathbf{x}^{*}\rangle$. Given the convexity of $f(\mathbf{x})$, we have $f(\mathbf{x})-f(\mathbf{x}^{*})\geq\langle\nabla f(\mathbf{x}^{*}),\mathbf{x}-\mathbf{x}^{*}\rangle\geq 0$, leading to $h(\mathbf{x})\geq 0$. As $h(\mathbf{x}^{*})=0$, this implies that $\mathbf{x}^{*}$ is the global minimum of $h$. Applying Lemma B.6, gives us,